OpenStudy (anonymous):

A bullet of 0.0500 kg is fired into a block of wood. Knowing that the bullet left the gun with a muzzle velocity of 350. m/s, and the bullet penetrates .15 m into the block of wood, determine: a) The average force required to stop the bullet. b) The impulse exerted by the wood on the bullet.

OpenStudy (anonymous):

question is this two objects colliding.....?

OpenStudy (anonymous):

Force=mass times acceleration

OpenStudy (anonymous):

ANSWER FOR A) 2.0x e4 N b) .17.5kg*m/s i tried f=ma didnt get the right answer...

OpenStudy (schrodingers_cat):

A. Time for the bullet to stop is d/vf = .15/(350 +0)/2 = 8.57 x 10^-4 \[\Delta p = \Sigma f \Delta t\] f = .05(350 -0)/ (8.57 X10 ^-4) = 2 x 10^4 N B. I = .05(350 -0) = 17.5 kg x m/s